The point (5, 7) lies on the line -2x + 8y = 46. Determine if this point is a solution of the system
-2x + 8y = 46
x – 2y = -9

Answer by robertb(5830)

It's enough to check if the coordinates of the point (5,7) satisfy the bottom equation x - 2y = -9:
5 - 2*7 = 5 - 14 = -9.
Hence the point is also on the line x - 2y = -9, and the point is a solution of the system.